OpenStudy (anonymous):

solve x+12/x+4=x/x+8

OpenStudy (unklerhaukus):

is that ; \[\frac{x+12}{x+4}=\frac x{x+8}\]

OpenStudy (whpalmer4):

Assuming it is, just cross-multiply, expand, and solve for \(x\): \[(x+12)(x+8) = (x+4)x\]

OpenStudy (anonymous):

In this special case (as |12-4|=|0-8|=8), subtract 1 from both sides and then cross-multiply. U will arrive at the answer quickly.